# Enriching Diagrams with Algebraic Operations

### 08 Dec 2023 v2

@@ Abstract
In this paper, we extend diagrammatic reasoning in monoidal categories with algebraic operations and equations. We achieve this by considering monoidal categories that are enriched in the category of Eilenberg-Moore algebras for a monad. Under the condition that this monad is monoidal and affine, we construct an adjunction between symmetric monoidal categories and symmetric monoidal categories enriched over algebras for the monad. This allows us to devise an extension, and its semantics, of the ZX-calculus with probabilistic choices by freely enriching over convex algebras, which are the algebras of the finite distribution monad. We show how this construction can be used for diagrammatic reasoning of noise in quantum systems.
@@ 1 Introduction
Monoidal categories are one way of generalizing algebraic reasoning and they can be used to draw intuitive diagrams that encapsulate this reasoning graphically. That monoidal categories are a powerful abstraction has been demonstrated in countless areas, such as linear logic [18] or quantum mechanics [1], just to name a few, and are amenable to graphical reasoning [43]. Another abstraction of algebraic reasoning are monads [3, 34, 40] and their algebras, or representations thereof [20, 33], which are distinct from monoidal categories in that identities (like associativity) always hold strictly and they allow rather arbitrary algebraic operations. In this paper, we set out to combine these two approaches into one framework, in which monoidal category diagrams can be composed not only sequentially and with a tensor product but also with additional algebraic operations.
@@
One such operation is the formation of convex combinations, which can be used to create a probabilistic mix of two or more diagrams. This occurs, for instance, when reasoning about the behaviour of noise in quantum circuits. Figure 1 shows on the left two gates, one called $G$ and one called $E$ that, respectively, model the wanted behaviour and a possible error. These two gates are mixed, where $G$ gets a probability of $0.9$ and $E$ of $0.1$. The trapezoids delimit the combination of the gates, and $A$ and $B$ are the input and output types of the gates [^1] We read diagrams from top to bottom.. In monoidal categories, the gates in the picture represent morphisms $G,E\colon A\to B$ and our aim is then to interpret the trapezoid block as a convex sum $G+_{0.9}E$ of these morphisms, where we define $G+_{p}E=pG+(1-p)E$. Such sums should also nicely interact with the tensor product. For instance, if $G_{1}\colon A\to B$ and $G_{2}\colon C\to D$ are gates, then an identity such as $(G_{1}+_{0.9}E)\otimes G_{2}=(G_{1}\otimes G_{2})+_{0.9}(E\otimes G_{2})$ should hold for these morphisms of type $A\otimes C\to B\otimes D$, see Figure 1 on the right. Having an operation to form convex combinations together with intuitive identities enables reasoning about e.g. probabilistic combination and noise in quantum circuits.
The difficulty lies then in combining monoidal diagrams with algebraic operations such that this interaction is coherent with both the monoidal identities and the algebraic identities. We will handle this difficulty by using enriched monoidal categories, where the enrichment yields the algebraic operations and the monoidal structure the parallel composition. More precisely, we will assume that the algebraic theory is given by a monad $T$ and that the monoidal categories are enriched over the Eilenberg-Moore category $\mathbf{Alg}^{T}$ of algebras for this monad. Our aim in this paper is to construct for an arbitrary monoidal category $\mathbf{C}$ an $\mathbf{Alg}^{T}$-enriched monoidal category $F\mathbf{C}$ that is free in the sense that there is an inclusion $\iota_{\mathbf{C}}\colon\mathbf{C}\to(F\mathbf{C})_{0}$ into the underlying category of $F\mathbf{C}$ and for every $\mathbf{Alg}^{T}$-enriched monoidal category $\underline{\mathbf{D}}$ and monoidal functor $G\colon\mathbf{C}\to\underline{\mathbf{D}}_{0}$, there is a unique $\mathbf{Alg}^{T}$-enriched monoidal functor $\bar{G}_{0}\colon F\mathbf{C}\to\underline{\mathbf{D}}$ that makes the following diagram commute.
@@
This free construction does not work for all monads, but we show that the free enrichment always exists for monoidal affine monads.
@@
Specifically, we contribute a definition of monoidal diagrams that are enriched with algebraic operations and a construction of free enrichment over algebras for monoidal affine monads. This enables composing diagrams in arbitrary monoidal categories with algebraic operations. We show how the theory can be applied to obtain convex combinations for ZX-diagrams and what the resulting identities of diagrams are. By exploiting the mapping property of the free enrichment, we obtain automatically sound interpretations of these operations and identities.
@@
ZX-diagrams are universal in the sense that they can technically represent any linear map between Hilbert spaces of dimension $\mathbb{C}^{2^{n}}$ [14]. Indeed, sums and linear/convex combinations [23, 45] of ZX-diagrams can be encoded within the language, but in practice these representations oftentimes lead to either very large diagrams or to diagrams that do not reveal upon visual inspection the (linear/convex) structure that the diagram is technically representing, which in return lessens the advantages gained by reasoning in terms of abstract graphical representations. Our perspective of using enrichment keeps the abstraction barrier and thus makes reasoning about convex combinations of diagrams tractable. In general, our theory covers also linear combinations such as [46, 36] and other, so far unexplored, algebraic operations such as those of join-semilattices. Moreover, the identities that have to be crafted carefully by hand and proven to be sound fall automatically out of our theory. Other related work is that of Sheet diagrams [15] and Tape diagrams [5], recently developed graphical languages for rig categories, which are categories with two monoidal structures – one for addition and one for multiplication.
@@ 2 Background
In this section we recall some needed definitions concerning category theory, for which we assume some (basic) familiarity, which otherwise we refer the reader to [32]. For brevity, we sometimes omit coherence conditions that some constructions have to satisfy.
We denote the collection of objects of a category C as $|\mathbf{C}|$, and the morphisms from object $A$ to $B$ as $\mathbf{C}(A,B)$. A monoidal category $(\mathbf{C},\otimes,\mathit{I})$ is a category C together with a functor $\otimes:\mathbf{C}\times\mathbf{C}\to\mathbf{C}$ called the tensor product and an object $\mathit{I}\in|\mathbf{C}|$ called the tensor unit subject to some conditions [26]. We will often refer to a monoidal category $(\mathbf{C},\otimes,\mathit{I})$ as just C. A monoidal category is a symmetric monoidal category (SMC) when it also has a braiding $\sigma_{A,B}:A\otimes B\to B\otimes A$ such that $\sigma_{B,A}\circ\sigma_{A,B}=\text{Id}_{A\otimes B}$.
@@
  • for all objects $A,B,C\in|\underline{\mathbf{C}}|$, a composition morphism ¯$\underline{\circ}\colon\underline{\mathbf{C}}(B,C)\times\underline{\mathbf{C}}(A,B)\to\underline{\mathbf{C}}(A,C)$ in $\mathbf{V}$, and
@@
subject to associativity and unit axioms [26]. We say that $\mathbf{V}$ is the base of enrichment for $\underline{\mathbf{C}}$. A way to look at the above definition is that we construct a $\mathbf{V}$-enriched category $\underline{\mathbf{C}}$ by identifying collections of morphisms of some category $\mathbf{C}$ as objects from $\mathbf{V}$, which we are able to compose by using the tensor product of $\mathbf{V}$. The most well-known example is that of locally small categories, in which the morphisms between two objects form a set, and thus we can see them as objects in the monoidal category ($\mathbf{Set},\times,\ast$) for $\times$ the Cartesian product and $\ast$ the singleton set.
With a suitable definition of $\mathbf{V}$-functors and $\mathbf{V}$-natural transformation, $\mathbf{V}$-categories organise themselves into a 2-category [26], that we denote by $\mathbf{V}\text{-}\mathbf{Cat}$. This category is also a symmetric monoidal category as follows. We define for $\mathbf{V}$-categories $\underline{\mathbf{C}}$ and $\underline{\mathbf{D}}$ a $\mathbf{V}$-category $\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}$ with objects $|\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}|=|\underline{\mathbf{C}}|\times|\underline{\mathbf{D}}|$ and hom-objects $(\underline{\mathbf{C}}\otimes\underline{\mathbf{D}})((A,B),(C,D))=\underline{\mathbf{C}}(A,C)\times\underline{\mathbf{D}}(B,D)$. The unit is given by $j_{(A,B)}=\ast\cong\ast\times\ast\xrightarrow{u_{A}\times v_{B}}\underline{\mathbf{C}}(A,A)\times\underline{\mathbf{D}}(B,B)$ in terms of the units $u$ of $\underline{\mathbf{C}}$ and $v$ of $\underline{\mathbf{D}}$. Similarly, one also defines the composition for $\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}$ in terms of the composition morphisms of $\underline{\mathbf{C}}$ and $\underline{\mathbf{D}}$, appealing to the symmetry in $\mathbf{V}$ [26, Sec. 1.4]. The tensor product also extends to $\mathbf{V}$-functors and $\mathbf{V}$-natural transformations, which makes it a 2-functor. Finally, one defines $I$ to be the unit $\mathbf{V}$-category with one object 00 and $I(0,0)=\ast$ and we thus obtain, with suitable definitions of associators etc., a symmetric monoidal 2-category $(\mathbf{V}\text{-}\mathbf{Cat},\otimes,I)$.
Most of the categories we are interested in are also dagger-compact categories (-$\dagger\text{-}CC$). These are SMC with some additional structure. First, they are equipped with an endofunctor $\dagger:\mathbf{C}^{op}\to\mathbf{C}$, that satisfies $(\text{Id}_{A})^{\dagger}=\text{Id}_{A}$, $(g\circ f)^{\dagger}=f^{\dagger}\circ g^{\dagger}$, $(f^{\dagger})^{\dagger}=f$, and $(f\otimes g)^{\dagger}=f^{\dagger}\otimes g^{\dagger}$. And secondly, for every object $A$ there exists a dual $A^{*}$ such that there exists unit $\eta_{A}:I\to A\otimes A^{*}$ and counit $\epsilon_{A}:A^{*}\otimes A\to I$ morphisms subject to some conditions [19].
@@ 3 Algebraic enrichment
The Distribution monad $\mathcal{D}$ is the functor in the category of sets $\mathcal{D}:\mathbf{Set}\to\mathbf{Set}$ that maps a set $A$ to the set $\mathcal{D}(A)$ of (finitely supported) probability distributions over elements of $A$. We write probability distributions as formal convex sums: $\sum_{a}p_{a}[a]\in\mathcal{D}(A)$ such that $a\in A,p_{a}\in[0,1]$, and $\sum_{a}p_{a}=1$. $\mathcal{D}$ acts on a morphism $f$ by simply applying $f$ to the underlying set: $(\mathcal{D}f)(\sum_{a}p_{a}[a])=\sum_{a}p_{a}[f(a)]$. The unit of $\mathcal{D}$ is the map $\eta:A\to\mathcal{D}(A):a\mapsto 1[a]$ (the Dirac distribution) and the multiplication $\mu$ of $\mathcal{D}$ “flattens” a distribution of distributions by multiplying the probabilities together: $\mu:\mathcal{D}(\mathcal{D}(A))\to\mathcal{D}(A):\sum_{q}p_{q}[\sum_{a}q_{a}[a]]\mapsto\sum_{a}r_{a}[a]$ where $r_{a}=\sum_{q}p_{q}q_{a}$ [22]. Moreover, the monad $\mathcal{D}$ is monoidal, meaning that there is a map:
@@
for every $A,B\in|\mathbf{Set}|$. It is also affine, meaning that $\mathcal{D}(\ast)\cong\ast$.
If $\mathcal{D}$ is a monad for expressing convex combinations of elements of a set, the Multiset monad $\mathcal{M}$ is its analogue for linear combinations with coefficients over some semiring $S$.
@@
When $T$ is a monoidal monad on a monoidal category $(\mathbf{C},\otimes,I)$, the tensor product of $T$ algebras $(A,a),(B,b)$, denoted $(A,a)\otimes_{T}(B,b)$, is (if it exists) defined as the coequalizer diagram [42, 9]
@@
making $(TA,\mu)\otimes_{T}(TB,\mu):=F(A\otimes B)$ [42, Prop. 2.5.2]. The monoidal unit is $T(I)$, while the associator, unitor, and symmetry (if present) are the images of the ones in $(\mathbf{C},\otimes,I)$ under $F$.
A functor $F:(\mathbf{V}_{1},\otimes_{\mathbf{V}_{1}},I_{\mathbf{V}_{1}})\to(\mathbf{V}_{2},\otimes_{\mathbf{V}_{2}},I_{\mathbf{V}_{2}})$ between two monoidal categories can be lifted to a $2$-functor $F_{*}:\mathbf{V}_{1}\text{-}\mathbf{Cat}\to\mathbf{V}_{2}\text{-}\mathbf{Cat}$ [7]. This is called a change of enriching, where we turn a $\mathbf{V}_{1}$-category into a $\mathbf{V}_{2}$-category. Indeed, given a $\mathbf{V}_{1}$-category $\underline{\mathbf{C}}$, we can construct the $\mathbf{V}_{2}$-category $F_{*}\underline{\mathbf{C}}$ by defining $|F_{*}\underline{\mathbf{C}}|:=|\underline{\mathbf{C}}|$ and, for every $A,B\in|F_{*}\underline{\mathbf{C}}|$, the hom-objects are $F_{*}\underline{\mathbf{C}}(A,B):=F(\underline{\mathbf{C}}(A,B))$ with composition and identity element following from the ones in $\underline{\mathbf{C}}$ under $F$.
@@
Let $(V,\otimes,I)$ be a closed symmetric monoidal category with coproducts. Then the forgetful functor $\mathbf{V}(I,-):\mathbf{V}\to\mathbf{Set}$ has a left adjoint $F$ that is a strong morphism of symmetric monoidal categories with natural isomorphisms $\mu,\epsilon$. Both functors induce a forgetful $2$-functor $\mathbf{V}(I,-)_{*}:\mathbf{V}\text{-}\mathbf{Cat}\to\mathbf{Cat}$ and a left adjoint $F_{*}$.
@@
A monoidal monad $T$ on $(\mathbf{Set},\times,\ast)$ endows the category of $T$-algebras $\mathbf{Alg}^{T}$ of bicomplete (complete and cocomplete) closed symmetric monoidal structure. This allows to lift the free-forgetful adjunction of Lemma 1 as a change of enriching between $\mathbf{Alg}^{T}$-categories and $\mathbf{Set}$-categories for a monoidal $T$.
@@
The proof follows from Lemma 1 and previous arguments. Given that $\mathbf{Alg}^{T}$ for $T$ a monad on $\mathbf{Set}$ is bicomplete, then coequalizer (1) exists and we can define tensor products of algebras. We can then make $\mathbf{Alg}^{T}$ a (symmetric) monoidal category given that $\mathbf{Set}$ is closed symmetric monoidal. Finally, $\mathbf{Alg}^{T}$ can be made into a closed category following [30] given that $\mathbf{Set}$ has equalizers. We can then use Lemma 1 to create a change of enriching between $\mathbf{Alg}^{T}$-categories and $\mathbf{Set}$-categories.
@@
Intuitively, we distinguish between probabilistic and deterministic processes by having the former enclosed within distribution brackets (in the same way as we would represent them as a formal sum $0.9[f]+0.1[g]$), that we choose to depict as trapezoids in this paper. Deterministic processes are depicted without the bracket enclosing mostly as syntactic sugar, otherwise they would simply have a single choice with probability $1$. We can see how wires can have weights inside this environment, and how each wire represents a probabilistic choice. Intuition also tells us that we could for example rewrite the diagram above by distributing $h$ over the two probabilistic branches. We will discuss in later sections which graphical rules capture the interactions present in these enriched categories. It is natural to then ask if our enriched category $\mathbf{C}$ maintained its monoidal structure, and if other desired properties (such as braiding and symmetry, if present) would still hold too. We will address this in the next section.

### 29 Jan 2024 v3

@@ Abstract
In this paper, we extend diagrammatic reasoning in monoidal categories with algebraic operations and equations. We achieve this by considering monoidal categories that are enriched in the category of Eilenberg-Moore algebras for a monad. Under the condition that this monad is monoidal and there is an adjunction between the free algebra functor and the underlying category functor, we construct an adjunction between symmetric monoidal categories and symmetric monoidal categories enriched over algebras for the monad. This allows us to devise an extension, and its semantics, of the ZX-calculus with probabilistic choices by freely enriching over convex algebras, which are the algebras of the finite distribution monad. We show how this construction can be used for diagrammatic reasoning of noise in quantum systems.
@@ 1 Introduction
Monoidal categories are one way of generalizing algebraic reasoning and they can be used to draw intuitive diagrams that encapsulate this reasoning graphically. That monoidal categories are a powerful abstraction has been demonstrated in countless areas, such as linear logic [19] or quantum mechanics [1], just to name a few, and are amenable to graphical reasoning [47] with diagrammatic languages such as the ZX-calculus [15]. Another abstraction of algebraic reasoning are monads [3, 37, 43] and their algebras, or representations thereof [21, 36], which are distinct from monoidal categories in that identities (like associativity) always hold strictly and they allow rather arbitrary algebraic operations. In this paper, we set out to combine these two approaches into one framework, in which monoidal category diagrams can be composed not only sequentially and in parallel with a tensor product but also with additional algebraic operations.
@@
One such operation is the formation of convex combinations, which can be used to create a probabilistic mix of two or more diagrams. This occurs, for instance, when reasoning about the behaviour of noise in quantum circuits. Figure 1 shows on the left two quantum logic gates, one called $G$ and one called $E$ that, respectively, model the wanted behaviour and a possible error. These two gates are mixed, where $G$ gets a probability of $0.9$ and $E$ of $0.1$. The trapezoids delimit the combination of the gates, and $A$ and $B$ are the input and output types of the gates [^1] We read diagrams from top to bottom.. In monoidal categories, the gates in the picture represent morphisms $G,E\colon A\to B$ and our aim is to interpret the trapezoid block as a convex sum $G+_{0.9}E$ of these morphisms, where we define $G+_{p}E=pG+(1-p)E$. Such sums should also nicely interact with the tensor product. For instance, if $G_{1}\colon A\to B$ and $G_{2}\colon C\to D$ are gates, then an identity such as $(G_{1}+_{0.9}E)\otimes G_{2}=(G_{1}\otimes G_{2})+_{0.9}(E\otimes G_{2})$ should hold for these morphisms of type $A\otimes C\to B\otimes D$, see Figure 1 on the right. Having an operation to form convex combinations together with intuitive identities enables reasoning about, for example, probabilistic combination and noise in quantum circuits.
The difficulty lies in combining monoidal diagrams with algebraic operations such that the algebraic identities and the monoidal identities interact coherently. We will handle this difficulty by using enriched monoidal categories, where the enrichment yields the algebraic operations and the monoidal structure the parallel composition. More precisely, we will assume that the algebraic theory is given by a monad $T$ and that the monoidal categories are enriched over the Eilenberg-Moore category $\mathbf{Alg}^{T}$ of algebras for this monad. Our aim in this paper is to construct for an arbitrary monoidal category $\mathbf{C}$ an $\mathbf{Alg}^{T}$-enriched monoidal category $F\mathbf{C}$ that is free in the sense that there is an inclusion $\iota_{\mathbf{C}}\colon\mathbf{C}\to(F\mathbf{C})_{0}$ into the underlying category of $F\mathbf{C}$ and for every $\mathbf{Alg}^{T}$-enriched monoidal category $\underline{\mathbf{D}}$ and monoidal functor $G\colon\mathbf{C}\to\underline{\mathbf{D}}_{0}$, there is a unique $\mathbf{Alg}^{T}$-enriched monoidal functor $\bar{G}_{0}\colon F\mathbf{C}\to\underline{\mathbf{D}}$ that makes the following diagram commute.
@@
This free construction does not work for all monads, but we show that the free enrichment always exists for monoidal $\mathbf{Set}$-monads whose free $T$-algebra functor is left adjoint to the underlying category functor $\mathbf{Alg}^{T}(I,-):\mathbf{Alg}^{T}\to\mathbf{Set}$ for $I$ the monoidal unit of $\mathbf{Alg}^{T}$.
@@
Specifically, we contribute a construction for free enrichment over algebras for some monoidal monads in Theorem 3.1 and Corollary 1. We also show how the enrichment preserves symmetric monoidal structure in Theorem 4.2 and Corollary 2. Given this construction, we demonstrate how a graphical language for reasoning in monoidal categories can be enriched with the free algebras for these monads, which enables diagrammatic reasoning of the interaction between the sequential and parallel compositions with the algebraic structure. We show how the theory can be applied to obtain convex combinations of ZX-diagrams and what the resulting identities of diagrams are. By exploiting the mapping property of the free enrichment, we automatically obtain sound interpretations of these operations and identities. Lastly, we describe how we can use the enrichment of ZX-diagrams to reason about noise in quantum systems.
@@
ZX-diagrams are universal in the sense that they can in principle represent any linear map between Hilbert spaces of dimension $\mathbb{C}^{2^{n}}$ [15]. Indeed, sums and linear/convex combinations [25, 49] of ZX-diagrams can be encoded within the language, but in practice these representations oftentimes lead to either very large diagrams or to diagrams that do not reveal upon visual inspection the (linear/convex) structure that the diagram is representing. This, in return, diminishes the advantages gained by reasoning in terms of abstract graphical representations. Our perspective of using enrichment keeps the abstraction barrier and thus makes reasoning about convex combinations of diagrams tractable. In general, our theory also covers the recently developed linear combinations of ZX-diagrams such as [50, 39] and other, so far unexplored, algebraic operations such as those of join-semilattices. Moreover, the identities that have to be crafted carefully by hand and proven to be sound fall automatically out of our theory. Other related work is that of Sheet diagrams [16] and Tape diagrams [5], recently developed graphical languages for rig categories, which are categories with two monoidal structures – one for addition and one for multiplication.
@@ 2 Background
In this section, we recall some terminology from category theory [8, 34, 35, 45] and introduce some notation. We denote the collection of objects of a category C as $|\mathbf{C}|$, and the morphisms from object $A$ to $B$ as $\mathbf{C}(A,B)$. A monoidal category $(\mathbf{C},\otimes,\mathit{I})$ is a category C together with a functor $\otimes:\mathbf{C}\times\mathbf{C}\to\mathbf{C}$ called the tensor product and an object $\mathit{I}\in|\mathbf{C}|$ called the tensor unit subject to some conditions [28]. We will often refer to a monoidal category $(\mathbf{C},\otimes,\mathit{I})$ as just C. A monoidal category is a symmetric monoidal category (SMC) when it also has a braiding $\sigma_{A,B}:A\otimes B\to B\otimes A$ such that $\sigma_{B,A}\circ\sigma_{A,B}=\text{Id}_{A\otimes B}$ that is also subject to coherence conditions [28].
@@
  • for objects $A,B,C\in|\underline{\mathbf{C}}|$, a composition morphism ¯$\underline{\circ}\colon\underline{\mathbf{C}}(B,C)\times\underline{\mathbf{C}}(A,B)\to\underline{\mathbf{C}}(A,C)$ in $\mathbf{V}$, and
@@
subject to associativity and unit axioms [28]. We say that $\mathbf{V}$ is the base of enrichment for $\underline{\mathbf{C}}$. A way to look at the above definition is that we construct a $\mathbf{V}$-enriched category $\underline{\mathbf{C}}$ by identifying morphisms of some category $\mathbf{C}$ as objects from $\mathbf{V}$, which we are able to compose by using the tensor product of $\mathbf{V}$. The most well-known example is that of locally small categories, in which the morphisms between two objects form a set, and thus we can see them as objects in the monoidal category ($\mathbf{Set},\times,\ast$) for $\times$ the Cartesian product and $\ast$ the singleton set.
With a suitable definition of $\mathbf{V}$-functors and $\mathbf{V}$-natural transformations, $\mathbf{V}$-categories organise themselves into a 2-category [28], denoted by $\mathbf{V}\text{-}\mathbf{Cat}$. For an SMC $(\mathbf{V},\times,\ast)$, $\mathbf{V}\text{-}\mathbf{Cat}$ is also an SMC as follows. We define for $\mathbf{V}$-categories $\underline{\mathbf{C}}$ and $\underline{\mathbf{D}}$ a $\mathbf{V}$-category $\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}$ with objects $|\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}|$ being the categorical product and hom-objects $(\underline{\mathbf{C}}\otimes\underline{\mathbf{D}})((A,B),(C,D))=\underline{\mathbf{C}}(A,C)\times\underline{\mathbf{D}}(B,D)$. The unit is given by $j_{(A,B)}=\ast\cong\ast\times\ast\xrightarrow{u_{A}\times v_{B}}\underline{\mathbf{C}}(A,A)\times\underline{\mathbf{D}}(B,B)$ in terms of the units $u$ of $\underline{\mathbf{C}}$ and $v$ of $\underline{\mathbf{D}}$. Similarly, one also defines the composition for $\underline{\mathbf{C}}\otimes\underline{\mathbf{D}}$ in terms of the composition morphisms of $\underline{\mathbf{C}}$ and $\underline{\mathbf{D}}$, appealing to the symmetry in $\mathbf{V}$ [28, Sec. 1.4]. The tensor product also extends to $\mathbf{V}$-functors and $\mathbf{V}$-natural transformations, which makes it a 2-functor. Finally, one defines $I$ to be the unit $\mathbf{V}$-category with one object 00 and $I(0,0)=\ast$ and we thus obtain, with suitable definitions of associators etc., a symmetric monoidal 2-category $(\mathbf{V}\text{-}\mathbf{Cat},\otimes,I)$.
Most of the categories we are interested in are also dagger-compact categories (-$\dagger\text{-}CC$). These are SMCs $(\mathbf{C},\otimes,\mathit{I})$ with some additional structure. First, they are equipped with an endofunctor $\dagger:\mathbf{C}^{op}\to\mathbf{C}$, that satisfies $(\text{Id}_{A})^{\dagger}=\text{Id}_{A}$, $(g\circ f)^{\dagger}=f^{\dagger}\circ g^{\dagger}$, $(f^{\dagger})^{\dagger}=f$, and $(f\otimes g)^{\dagger}=f^{\dagger}\otimes g^{\dagger}$. And secondly, for every object $A$ there exists a dual $A^{*}$ such that there exists unit $\eta_{A}:I\to A\otimes A^{*}$ and counit $\epsilon_{A}:A^{*}\otimes A\to I$ morphisms subject to some conditions [20].
@@ 3 Algebraic Enrichment
The Distribution monad $(\mathcal{D},\mu,\eta)$ [^2] We will often refer to a monad $(T,\mu,\eta)$ as $T$. contains the functor $\mathcal{D}:\mathbf{Set}\to\mathbf{Set}$ that maps a set $A$ to the set $\mathcal{D}(A)$ of (finitely supported) probability distributions over elements of $A$. We write probability distributions as formal convex sums: $\sum_{a}p_{a}[a]\in\mathcal{D}(A)$ such that $a\in A,p_{a}\in[0,1]$, and $\sum_{a}p_{a}=1$. $\mathcal{D}$ acts on a morphism $f$ by simply applying $f$ to the underlying set: $(\mathcal{D}f)(\sum_{a}p_{a}[a])=\sum_{a}p_{a}[f(a)]$. The unit of the monad is the map $\eta:A\to\mathcal{D}(A):a\mapsto 1[a]$ (the Dirac distribution), and the multiplication $\mu$ “flattens” a distribution of distributions by multiplying the probabilities together: $\mu:\mathcal{D}(\mathcal{D}(A))\to\mathcal{D}(A):\sum_{q}p_{q}[\sum_{a}q_{a}[a]]\mapsto\sum_{a}r_{a}[a]$ where $r_{a}=\sum_{q}p_{q}q_{a}$ [23].
The functor $\mathcal{D}$ is also a monoidal functor, which makes $(\mathcal{D},\mu,\eta)$ a monoidal monad. In particular, this means that there exists a map:
@@
for every $A,B\in|\mathbf{Set}|$.
A monad $T:\mathbf{C}\to\mathbf{C}$ is affine if there is an isomorphism $T(\ast)\cong\ast$ for $\ast$ the terminal object of $\mathbf{C}$ [24]. This is the case for $\mathcal{D}$.
If $\mathcal{D}$ is a monad for expressing convex combinations of elements of a set, the Multiset monad $\mathcal{M}$ is its analogue for linear combinations with coefficients over some semiring.
@@
When $T$ is a monoidal monad on a monoidal category $(\mathbf{C},\otimes,I)$, the tensor product of $T$ algebras $(A,a),(B,b)$, denoted $(A,a)\otimes^{T}(B,b)$, is (if it exists) defined as the coequalizer diagram [46, 10]
@@
Therefore, $(T(A),\mu)\otimes^{T}(T(B),\mu)\cong F(A\otimes B)$ [46, Prop. 2.5.2]. The monoidal unit is $I^{T}=(T(I),\mu)$, while the associator, unitor, and symmetry (if present) are the images of the ones in $(\mathbf{C},\otimes,I)$ under $F$. We then have that $(\mathbf{Alg}^{T},\otimes^{T},I^{T})$ is the (symmetric) monoidal category of free $T$-algebras.
A functor $F:(\mathbf{V}_{1},\otimes_{\mathbf{V}_{1}},I_{\mathbf{V}_{1}})\to(\mathbf{V}_{2},\otimes_{\mathbf{V}_{2}},I_{\mathbf{V}_{2}})$ between two monoidal categories can be lifted to a $2$-functor $F_{*}:\mathbf{V}_{1}\text{-}\mathbf{Cat}\to\mathbf{V}_{2}\text{-}\mathbf{Cat}$ [7]. This is called a change of enriching, where we turn a $\mathbf{V}_{1}$-category into a $\mathbf{V}_{2}$-category. Indeed, given a $\mathbf{V}_{1}$-category $\underline{\mathbf{C}}$, we can construct the $\mathbf{V}_{2}$-category $F_{*}\underline{\mathbf{C}}$ by defining $|F_{*}\underline{\mathbf{C}}|:=|\underline{\mathbf{C}}|$ and, for every $A,B\in|F_{*}\underline{\mathbf{C}}|$, the hom-objects are $F_{*}\underline{\mathbf{C}}(A,B):=F(\underline{\mathbf{C}}(A,B))$ with composition and identity element following from the ones in $\underline{\mathbf{C}}$ under $F$. For a symmetric monoidal category $(\mathbf{V},\otimes,I)$, an important instance is the functor $\mathbf{V}(I,-)_{*}\colon\mathbf{V}\text{-}\mathbf{Cat}\to\mathbf{Cat}$ called the underlying category functor and it is denoted by $(-)_{0}$.
@@
Let $(\mathbf{V},\otimes,I)$ be a closed symmetric monoidal category with coproducts. Then the hom-functor $\mathbf{V}(I,-)\colon\mathbf{V}\to\mathbf{Set}$ has a left adjoint $F$ that sends a set $X$ to $F(X)=\coprod_{X}I$, the $X$-th fold copower of $I$. Moreover, $F$ is a strong morphism of symmetric monoidal categories and the induced $2$-functor $F_{*}$ is left-adjoint to the underlying category functor $(-)_{0}$.
@@
A monoidal monad $T$ on $(\mathbf{Set},\times,\ast)$ endows the category of $T$-algebras $\mathbf{Alg}^{T}$ with a bicomplete (complete and cocomplete) closed SMC structure. This allows to lift the free-forgetful adjunction of Lemma 1 as a change of enriching between $\mathbf{Alg}^{T}$-categories and $\mathbf{Set}$-categories for a monoidal $T$.
@@
The proof follows from Lemma 1 and previous arguments. Given that $\mathbf{Alg}^{T}$ for $T$ a monad on $\mathbf{Set}$ is bicomplete, then coequalizer (1) exists and we can define tensor products of algebras. We can then make $\mathbf{Alg}^{T}$ a symmetric monoidal category given that $\mathbf{Set}$ is closed symmetric monoidal. Finally, $\mathbf{Alg}^{T}$ can be made into a closed category following [32] given that $\mathbf{Set}$ has equalizers. We can then use Lemma 1 to create a change of enriching between $\mathbf{Alg}^{T}$-categories and $\mathbf{Set}$-categories.
Corollary 1
Let $T$ be a monoidal monad on $\mathbf{Set}$ defined by a free-forgetful adjunction $U:\mathbf{Alg}^{T}\rightleftarrows\mathbf{Set}:L$. If $L$ is naturally isomorphic to the functor $F$ from Lemma 1, that is $(T(-),\mu)\cong\coprod_{(-)}(TI,\mu)$, then the induced $2$-functor $L_{*}$ is left adjoint to the underlying category functor $(-)_{0}$. This lets us use Theorem 3.1 to enrich locally small categories with the free algebras over $T$. The condition $L\cong F$ holds, in particular, when $T$ is an affine monad.
Proof
Whenever we have that $L\cong F$, the enrichment over free $T$-algebras comes simply from substituting $F$ with $L$ in Lemma 1 and Theorem 3.1. To see that this condition holds when $T$ is affine, we construct hom isomorphisms $\mathbf{Alg}^{T}(L(X),Y)\cong\mathbf{Set}(X,U(Y))\cong\mathbf{Set}(X,\mathbf{Alg}^{T}(I^{T},Y))\cong\mathbf{Alg}^{T}(F(X),Y)$ for some $X\in|\mathbf{Set}|,Y\in|\mathbf{Alg}^{T}|$, with the second and last isomorphism coming from their respective adjunctions. The remaining one is due to $T(*)\cong*$, which allows us to get algebra homomorphisms $h:(T(*),\mu)\mapsto(T(Y),\mu)$ from maps $h^{\prime}:*\mapsto Y$, while the other direction just requires to forget the homomorphism structure.
@@
Intuitively, we distinguish between probabilistic and deterministic processes by having the former enclosed within distribution brackets (in the same way as we would represent them as a formal sum $0.9[f]+0.1[g]$), that we choose to depict as trapezoids in this paper. Deterministic processes are depicted without the bracket enclosing mostly as syntactic sugar, otherwise they would simply have a single choice with probability $1$. We can see how wires can have weights inside this environment, and how each wire represents a probabilistic choice. Intuition also tells us that we could for example rewrite the diagram above by distributing $h$ over the two probabilistic branches. We will discuss in later sections which graphical rules capture the interactions present in these enriched categories.
It is natural to then ask if our enriched category $\mathbf{C}$ maintained its monoidal structure, and if other desired properties (such as braiding and symmetry, if present) would still hold too. We will address this in the next section.